# ZeroDVFS: Zero-Shot LLM-Guided Core and Frequency Allocation for Embedded Platforms

### v1

@@ Appendix A LLM-Based Feature Extraction: Supplementary Details
This appendix provides comprehensive implementation details for the LLM-based semantic feature extraction methodology described in Section 4.4. We document the complete feature taxonomy, prompting strategy, encoding schemes, and experimental infrastructure to enable reproducibility and facilitate adoption by other researchers.
A.1 Semantic Feature Taxonomy
@@
Table 9 presents the complete taxonomy of 13 semantic features extracted through LLM analysis, organized into three categories that reflect distinct aspects of workload characterization relevant to DVFS scheduling decisions.
@@
A.2 Prompting Strategy and Implementation
We employ a zero-shot prompting approach with strict output formatting to ensure consistent feature extraction across all three LLMs. Figure 10 presents the condensed prompt template.
@@ A.2 Prompting Strategy and Implementation
A.3 Feature Encoding for Machine Learning
Categorical features require numerical encoding for integration with gradient boosting models. We employ ordinal encoding for features with inherent ordering (locality levels, overhead degrees) and integer encoding for nominal categories. Table 10 summarizes the encoding mappings.
@@ A.3 Feature Encoding for Machine Learning
A.4 Experimental Dataset
The evaluation dataset comprises 47,040 total samples across two embedded platforms, generated by executing 42 OpenMP benchmarks across all frequency-core configurations. Table 11 summarizes the dataset composition.
@@ A.4 Experimental Dataset
A.5 Inter-Model Agreement Analysis
To assess the reliability of LLM-extracted features, we analyze agreement rates across the three models. Table 12 presents pairwise and unanimous agreement percentages for each semantic feature.
@@ A.5 Inter-Model Agreement Analysis
A.6 Computational Cost and Latency
Table 13 presents a detailed breakdown of computational costs and latency for the LLM feature extraction pipeline.
@@ A.6 Computational Cost and Latency
All experiments were conducted in December 2024 using official API endpoints: DeepSeek-V3 (deepseek-chat), Claude Sonnet (claude-3-5-sonnet), and GPT-4o (gpt-4o-2024-08-06). Commercial LLM outputs may vary over time due to model updates. For fully reproducible research, open-source alternatives such as DeepSeek-Coder, CodeLlama, or Qwen2.5-Coder can be deployed locally.
Table 14 compares end-to-end latency against traditional exhaustive profiling, demonstrating speedups exceeding three orders of magnitude.
@@
A.7 Cost Projections at Scale
Table 15 projects costs for large-scale deployments, comparing LLM extraction against traditional manual profiling.
@@ A.8 Robustness Analysis
For production deployments, we recommend ensemble voting across multiple LLMs for critical features, with disagreement flagging uncertain extractions for review. Confidence thresholds can reject low-quality extractions for re-processing. Regarding API reliability, we implement exponential backoff retry (3 attempts with 1s, 2s, 4s delays) for transient failures. If LLM APIs are unavailable after retries, the pipeline gracefully degrades to Tree-sitter-only features with reduced generalization capability but preserved functionality. Since LLM extraction is one-time per benchmark (features are cached), temporary API unavailability does not affect subsequent scheduling decisions which rely only on the 358ms RL inference (see latency discussion in Section LABEL:sec:latency_analysis).
A.9 Error Analysis by Benchmark Category
Table 16 presents prediction accuracy stratified by benchmark category, revealing systematic patterns in model performance.
@@ A.9 Error Analysis by Benchmark Category
A.10 Limitations and Future Directions

### v2

@@ Appendix B LLM-Based Feature Extraction: Supplementary Details
This appendix provides comprehensive implementation details for the LLM-based semantic feature extraction methodology described in the main paper’s LLM feature extraction section. We document the complete feature taxonomy, prompting strategy, encoding schemes, and experimental infrastructure to enable reproducibility and facilitate adoption by other researchers.
B.1 Motivation: Why LLMs for Workload Characterization?
B.2 Semantic Feature Taxonomy
@@
The table below presents the complete taxonomy of 13 semantic features extracted through LLM analysis, organized into three categories that reflect distinct aspects of workload characterization relevant to DVFS scheduling decisions.
@@
B.3 Prompting Strategy and Implementation
We employ a zero-shot prompting approach with strict output formatting to ensure consistent feature extraction across all three LLMs. The listing below presents the condensed prompt template.
@@ B.3 Prompting Strategy and Implementation
B.4 Feature Encoding for Machine Learning
Categorical features require numerical encoding for integration with gradient boosting models. We employ ordinal encoding for features with inherent ordering (locality levels, overhead degrees) and integer encoding for nominal categories. The table below summarizes the encoding mappings.
@@ B.4 Feature Encoding for Machine Learning
B.5 Experimental Dataset
The evaluation dataset comprises 47,040 total samples across two embedded platforms, generated by executing 42 OpenMP benchmarks across all frequency-core configurations. The table below summarizes the dataset composition.
@@ B.5 Experimental Dataset
B.6 Inter-Model Agreement Analysis
To assess the reliability of LLM-extracted features, we analyze agreement rates across the three models. The table below presents pairwise and unanimous agreement percentages for each semantic feature.
@@ B.6 Inter-Model Agreement Analysis
B.7 Feature Importance Ranking
The table below presents the complete feature importance ranking extracted from the XGBoost prediction model using the gain metric. This table was moved from the main text (originally Table 7) to save space. The ranking reveals the relative contribution of each feature to execution time prediction accuracy.
Hardware runtime metrics occupy the top 9 positions, collectively accounting for over 80% of total predictive importance. Energy measurements from various subsystems (main CPU, system, GPU, DDR, Denver cores) dominate because they have direct causal relationships with execution time. Programs consuming more energy generally run longer. Context switches rank second (22.3%), capturing OS scheduling overhead and task migration costs. Instruction counts and cache-related metrics (references and misses) capture computational complexity and memory hierarchy efficiency.
LLM-extracted semantic features appear starting at rank 10. Among the 13 semantic features, only two appear in the top 15: cache behavior pattern (rank 10, 1.03%) and data dependency type (rank 15, 0.57%). These features provide modest but meaningful contributions in the presence of hardware counters. However, their true value emerges in zero-shot prediction scenarios where hardware counters are unavailable. In such cases, semantic features become the primary predictive signal for new, unprofiled programs.
Table 13: Top 15 Features by Importance (Claude + TX2 + All Features Configuration)
Rank Feature Name Importance (%)
1 energy_main_j 28.52
2 context_switches 22.27
3 energy_system_j 16.59
4 instructions 7.15
5 energy_gpu_j 4.80
6 energy_ddr_j 3.23
7 cache_references 2.72
8 energy_denver_j 2.22
9 energy_cpu_j 1.41
10 cache_behavior_pattern (LLM) 1.03
11 num_cores 0.75
12 cycles 0.59
13 cache_misses 0.59
14 power_cpu_w 0.57
15 data_dependency_type (LLM) 0.57
  • Shaded rows indicate LLM-extracted semantic features.
  • Importance computed using XGBoost’s gain metric (proportion of total gain contributed by each feature).
  • Remaining 37 features (not shown) contribute the remaining  10% of importance.
The aggregate contribution of all 13 LLM features totals approximately 5% of total importance. While this may seem modest, it reflects the experimental scenario where hardware counters are available. In such scenarios, semantic features provide redundant information that the model appropriately down-weights. The critical distinction is that semantic features are predictive (describing code behavior before execution), whereas hardware counters are observational (measuring behavior during execution). This makes semantic features indispensable for zero-shot generalization to new programs.
B.8 Computational Cost and Latency
The table below presents a detailed breakdown of computational costs and latency for the LLM feature extraction pipeline.
@@ B.8 Computational Cost and Latency
All experiments were conducted in December 2024 using official API endpoints: DeepSeek-V3 (deepseek-chat), Claude Sonnet (claude-3-5-sonnet), and GPT-4o (model ID: gpt-4o-2024-
08-06
). Commercial LLM outputs may vary over time due to model updates. For fully reproducible research, open-source alternatives such as DeepSeek-Coder, CodeLlama, or Qwen2.5-Coder can be deployed locally.
The table below compares end-to-end latency against traditional exhaustive profiling, demonstrating speedups exceeding three orders of magnitude.
@@
The table below provides a comprehensive comparison of LLM-based feature extraction versus traditional exhaustive profiling across key practical deployment dimensions.
Table 16: LLM Feature Extraction vs. Traditional Profiling: Comprehensive Comparison
Metric LLM Extraction Traditional
Time per program <<5 seconds 8 to 12 hours
Monetary cost per program \$0.0015 to \$0.018 Hardware + electricity
Requires program execution No Yes
Requires target hardware No Yes
Generalizes to new programs Yes No
Cross-platform applicability Yes Platform-specific
This comparison demonstrates the fundamental advantages of LLM-based semantic feature extraction for zero-shot workload characterization. The elimination of execution and hardware requirements enables deployment scenarios infeasible with traditional profiling approaches.
B.9 Cost Projections at Scale
The table below projects costs for large-scale deployments, comparing LLM extraction against traditional manual profiling.
@@ B.10 Robustness Analysis
For production deployments, we recommend ensemble voting across multiple LLMs for critical features, with disagreement flagging uncertain extractions for review. Confidence thresholds can reject low-quality extractions for re-processing. Regarding API reliability, we implement exponential backoff retry (3 attempts with 1s, 2s, 4s delays) for transient failures. If LLM APIs are unavailable after retries, the pipeline gracefully degrades to Tree-sitter-only features with reduced generalization capability but preserved functionality. Since LLM extraction is one-time per benchmark (features are cached), temporary API unavailability does not affect subsequent scheduling decisions which rely only on the 358ms RL inference (see latency discussion in the main paper’s experiments section and the Decision Latency subsection of this appendix).
B.11 Error Analysis by Benchmark Category
The table below presents prediction accuracy stratified by benchmark category, revealing systematic patterns in model performance.
@@ B.11 Error Analysis by Benchmark Category
B.12 Limitations and Future Directions
@@
Appendix C Addressing Real-Time Systems Considerations
This section explicitly addresses key concerns regarding the applicability and limitations of ZeroDVFS within the context of real-time systems, directly responding to anticipated reviewer questions.
C.1 Scope Clarification: Average-Case vs. Worst-Case Optimization
System Classification. ZeroDVFS is designed for soft real-time and best-effort workloads where average-case performance optimization and energy efficiency are paramount, rather than hard real-time systems requiring formal worst-case execution time (WCET) guarantees and deadline proofs. The system does not provide formal WCET bounds or deadline guarantees, schedulability analysis under worst-case conditions, timing correctness proofs for safety-critical tasks, or protection against all possible deadline violations.
Intended Use Cases. ZeroDVFS targets applications where tasks have soft deadlines (occasional misses tolerable), energy efficiency and thermal management are critical optimization objectives, workload phases change dynamically requiring runtime adaptation, and average-case performance is more important than worst-case guarantees. Representative applications include mobile video processing, IoT sensor fusion, background analytics, and adaptive vision pipelines.
Inappropriate Use Cases. ZeroDVFS should not be deployed in safety-critical hard real-time systems (automotive brake controllers, medical devices, avionics), mixed-criticality systems where high-criticality tasks require WCET guarantees, applications governed by real-time standards (AUTOSAR, DO-178C) requiring formal certification, or time-triggered architectures requiring deterministic table-driven scheduling.
C.2 Fail-Safe Thermal Protection Mechanisms
Given that the environment model can exhibit high prediction errors during initial zero-shot deployment (64–73% MAPE on new platforms), the system implements multiple layers of thermal protection to prevent hardware damage during the adaptation phase. During the initial zero-shot deployment period (first 10–20 episodes on a new platform), before the model has been fine-tuned with platform-specific samples, the system operates under conservative constraints. The maximum frequency is limited to 50% of the hardware-supported maximum, the maximum active cores are limited to 50% of available cores, the thermal reward penalty threshold is reduced from 50°C to 40°C, and epsilon-greedy exploration is disabled ($\epsilon=0$), using only greedy exploitation. This conservative mode ensures that even if the environment model wildly overestimates performance (predicting low temperature when actual temperature would be high), the physical frequency cap prevents reaching critical thermal zones.
The system also tracks prediction uncertainty using ensemble variance across multiple forward passes with dropout. Before applying each predicted action, the system runs 10 forward passes through the environment model with dropout enabled and computes prediction variance. If $\sigma_{\text{pred}}>\tau_{\text{threshold}}$ (e.g., $\tau=0.15$), the prediction is rejected, the system falls back to the Linux ondemand governor for that scheduling decision, and the rejected prediction is logged for offline analysis. This mechanism prevents the system from acting on highly uncertain predictions during the high-error adaptation phase.
Independent of the RL agent’s decisions, a hardware monitoring daemon runs continuously, sampling thermal sensors every 100ms via sysfs (path: /sys/devices/virtual/
thermal/thermal_zone*/temp
). If any core exceeds 60°C (10°C below TX2’s 70°C throttling threshold), all cores are immediately reverted to minimum frequency. If temperature exceeds 65°C, the system triggers emergency fallback to the Linux powersave governor and logs a critical thermal event. Temperature must drop below 50°C before RL control is re-enabled. This hardware-level watchdog provides guaranteed thermal protection regardless of RL prediction accuracy.
As the model collects fine-tuning samples and prediction accuracy improves, constraints are gradually relaxed. After 5 samples, the frequency cap increases to 65% of maximum; after 10 samples, to 80% of maximum; and after 20 samples, all conservative constraints are removed and full RL control is enabled. Progression is tracked via validation MAPE: if MAPE << 50%, the system proceeds to the next relaxation stage. This phased approach balances thermal safety during high-error periods with eventual full-performance operation once the model adapts.
C.3 Runtime Decision Latency and Applicability Constraints
The Python-based prototype implementation exhibits profiler model inference latency of 122ms, thermal model inference latency of 122ms, and policy network inference latency of 122ms (including both Profiler and Temperature agents), resulting in a total per-decision overhead of 358ms. Critically, all reported experimental results (Tables 4, 5, Figure 6) include this 358ms overhead in the end-to-end makespan measurements. The measured 1.13s execution time for BOTS FFT includes the RL decision time as part of the total workload completion time. This ensures the evaluation reflects real-world deployment performance rather than hypothetical best-case scenarios.
Given the 358ms decision latency, ZeroDVFS is currently applicable only to coarse-grained workloads with minimum task duration of 1–5 seconds (where overhead represents 7–36% of execution time). The system is inappropriate for fine-grained control loops at 100Hz–1000Hz (10ms–1ms periods), requiring 358ms to react to workload phase transitions. For high-frequency embedded control applications (drone flight controllers, automotive trajectory planning, robotic servo loops), the current implementation’s 358ms latency is orders of magnitude too slow. By the time the RL agent computes an action, the control deadline has already expired.
Production deployment improvements outlined in future work include C++ rewrite to eliminate Python interpreter overhead, TensorRT optimization with FP16 quantization, kernel fusion, and INT8 calibration, projected to achieve sub-10ms latency based on measured TensorRT FP16 speedups on Jetson platforms (typical 20–40× acceleration over Python+TensorFlow). This would enable applicability to control loops at 10–100Hz (100ms–10ms periods). Importantly, the current paper is evaluated on the Python prototype. Production deployment claims are speculative future work, not empirical results. An ECRTS Tools & Implementations track paper must be rigorously evaluated on the actual implemented artifact.
C.4 Response to Baseline Equity Concerns
The paper compares ZeroDVFS against the Precise Scheduler [7], a hard real-time federated DAG scheduler designed for Mixed-Criticality Systems. The comparison is inherently asymmetrical and must be interpreted carefully. The Precise Scheduler guarantees formal worst-case timing correctness for all tasks under any execution scenario, deadline satisfaction even under maximum microarchitectural interference (cache thrashing, branch mispredictions, memory contention), mathematically provable schedulability analysis, and safety for mixed-criticality workloads where deadline misses are unacceptable. In contrast, ZeroDVFS optimizes average-case energy consumption and makespan, provides dynamic runtime adaptation to observed workload characteristics, implements thermal management through reactive frequency scaling, and provides no formal guarantees against deadline violations.
The Precise Scheduler consumes 75.5mJ versus ZeroDVFS’s 9.1mJ, representing a 66.4mJ delta. This energy difference is not an algorithmic inefficiency. It represents the worst-case safety margin required to maintain higher voltage/frequency to ensure deadline satisfaction under worst-case execution paths, thermal headroom to prevent thermal throttling during worst-case power spikes, schedulability slack from conservative core allocation to guarantee timing correctness, and the physical cost of predictability. If ZeroDVFS encounters a sudden worst-case execution path (e.g., 100% L2 cache miss rate, maximum branch mispredictions), its dynamically lowered frequency may cause a deadline miss that the Precise Scheduler would prevent by maintaining conservative margins.
Despite the asymmetry, the comparison quantifies the energy-predictability tradeoff. For soft real-time applications tolerating occasional misses, ZeroDVFS offers 7× energy savings. For hard real-time systems requiring formal guarantees, the Precise Scheduler’s 7× energy cost is justified. System designers can make informed tradeoffs based on application criticality. The comparison does not claim ZeroDVFS is ”better”. Rather, it demonstrates that average-case optimization and worst-case guarantees occupy different points on the Pareto frontier.
C.5 Deadline-Awareness and Timing Requirements
ZeroDVFS’s reward functions optimize for $r_{\text{profiler}}=f(\text{energy},\text{makespan})$ and $r_{\text{temp}}=g(\text{temperature})$. Notably absent are task deadlines, period constraints, or criticality levels. The system has no mechanism to prioritize high-criticality tasks over low-criticality tasks, guarantee that a task completes before its absolute deadline, adjust frequency scaling based on remaining time until deadline, or perform admission control or schedulability tests. If the LLM hallucinates and incorrectly characterizes a memory-bound workload as compute-bound, the environment model may predict high performance at low frequency. The RL agent, trusting this prediction, assigns insufficient frequency, causing the task to miss its deadline. The system has no deadline-aware recovery mechanism.
Deadline-awareness could be integrated via reward augmentation in future work:
$r_{\text{profiler}}^{\text{deadline}}=\alpha\cdot f(\text{energy},\text{makespan})+\beta\cdot h(\text{deadline\_slack})$
where $\text{deadline\_slack}=\text{deadline}-\text{predicted\_completion\_time}$ and $h(\cdot)$ applies exponentially increasing penalties as slack approaches zero. This would require annotating benchmarks with deadlines and periods, extending the environment model to predict completion time distributions (not just point estimates), integrating worst-case prediction intervals for conservative scheduling, and validation against hard real-time workloads (automotive, avionics benchmarks). This extension is explicitly listed as future work in the conclusion. The current system deliberately focuses on deadline-agnostic energy-thermal optimization.